ADNAN MUHAMMED

Grade 12,

extend the conductivity expression in terms of mobility and explain how carrier concentration and mobility vary with respect to temperature? plz reply someone as soon as possible..... thanku.....

extend the conductivity expression in terms of mobility and explain how carrier concentration and mobility vary with respect to temperature?
plz reply someone as soon as possible.....
thanku.....

Grade:12th pass

1 Answers

Yash Jain
9 years ago
I suppose this question is from Semiconductors
See, for that we have ITotal = Ie + Ih
If { V is the potential difference across the semiconductor,
R is the resistance through the semiconductor,
e is the charge of electron and hole, 
Ne and Nh are the respective carrier densities,
A is the area of cross section of the semiconductor,
ve and vh are the drift velocities of electrons and holes respectively }
then, V/R = NeeAve + NheAvh 
E.l / pl/A = eA.(Neve + Nhvh)
therefore, 1/p = e(N
eve/E + Nhvh/E)
or 1/p = S = e(Neue + Nhuh)
where p is the resistivity, S is conductivity and ue,uh are respective mobilities.
As the temperature increases, the mobility of the carriers decrease upto certain extend but there is large increment in their concentrations. So there is a considerable increase in the conductivity of semiconductors wrt the conductivity expression.
Hope that’ll help.